OpenStudy (anonymous):

step by step solve 4/3x + 3 =-5

OpenStudy (anonymous):

So, first add the reciprocal of 4/3x to each side? Then add 3 to each side?

OpenStudy (cwtan):

-3 for each side first then multiply by 3x

OpenStudy (anonymous):

ok this is what I did...-3/4x x 4/-3x=1x, -5 + 3 =8, the 1x +3 =4x/8=2x

OpenStudy (anonymous):

ok, then 4/3x +3-3=-5-3=4/3x x 3/4=3s cancel out; 4/2=2, right?

OpenStudy (noseboy908):

^is the work written out

OpenStudy (noseboy908):

Did that answer your question?